Linear Mass Balances

Linear Mass Balances – p.1/??

The buck stops here

Anonymous

(sign on President Truman’s desk)

Linear Mass Balances – p.2/??

Solution of Mass Balance Equations

The linear mass balances on various units are combined to analyze the ethanol process.

Step 1: Guess P and T levels in the flowsheet. Specify recoveries, split fractions, ... (use degrees of freedom for each unit).

Step 2: Determine coefficients for linear models in each unit (α_{k/n}, β, N, ξ).

Step 3: Setup linear equations and solve for flowrates of each component.

Step 4: Check guessed values from the first step. Calculate P and T from flowrates. If different from Step 1, go to Step 2 using these values of P and T. If flowsheet does not meet specifications, change P and/or T or modify flowsheet.

Linear Mass Balances – p.3/??

The linear mass balances on various units are combined to analyze the ethanol process.

Step 1: Guess P and T levels in the flowsheet. Specify recoveries, split fractions, ... (use degrees of freedom for each unit).

Step 2: Determine coefficients for linear models in each unit (α_{k/n}, β, N, ξ).

Step 3: Setup linear equations and solve for flowrates of each component.

Step 4: Check guessed values from the first step. Calculate P and T from flowrates. If different from Step 1, go to Step 2 using these values of P and T. If flowsheet does not meet specifications, change P and/or T or modify flowsheet.

Linear Mass Balances – p.3/??

The linear mass balances on various units are combined to analyze the ethanol process.

Step 1: Guess P and T levels in the flowsheet. Specify recoveries, split fractions, ... (use degrees of freedom for each unit).

Step 2: Determine coefficients for linear models in each unit (α_{k/n}, β, N, ξ).

Step 3: Setup linear equations and solve for flowrates of each component.

Step 4: Check guessed values from the first step. Calculate P and T from flowrates. If different from Step 1, go to Step 2 using these values of P and T. If flowsheet does not meet specifications, change P and/or T or modify flowsheet.

Linear Mass Balances – p.3/??

The linear mass balances on various units are combined to analyze the ethanol process.

Step 1: Guess P and T levels in the flowsheet. Specify recoveries, split fractions, ... (use degrees of freedom for each unit).

Step 2: Determine coefficients for linear models in each unit (α_{k/n}, β, N, ξ).

Step 3: Setup linear equations and solve for flowrates of each component.

Step 4: Check guessed values from the first step. Calculate P and T from flowrates. If different from Step 1, go to Step 2 using these values of P and T. If flowsheet does not meet specifications, change P and/or T or modify flowsheet.

Linear Mass Balances – p.3/??

The linear mass balances on various units are combined to analyze the ethanol process.

Step 1: Guess P and T levels in the flowsheet. Specify recoveries, split fractions, ... (use degrees of freedom for each unit).

Step 2: Determine coefficients for linear models in each unit (α_{k/n}, β, N, ξ).

Step 3: Setup linear equations and solve for flowrates of each component.

Step 4: Check guessed values from the first step. Calculate P and T from flowrates. If different from Step 1, go to Step 2 using these values of P and T. If flowsheet does not meet specifications, change P and/or T or modify flowsheet.

Linear Mass Balances – p.3/??

The linear mass balances on various units are combined to analyze the ethanol process.

Step 1: Guess P and T levels in the flowsheet. Specify recoveries, split fractions, ... (use degrees of freedom for each unit).

Step 2: Determine coefficients for linear models in each unit (α_{k/n}, β, N, ξ).

Step 3: Setup linear equations and solve for flowrates of each component.

Step 4: Check guessed values from the first step. Calculate P and T from flowrates. If different from Step 1, go to Step 2 using these values of P and T. If flowsheet does not meet specifications, change P and/or T or modify flowsheet.

Linear Mass Balances – p.3/??